[The following information applies to the questions displayed below.]

Nabisco Company's balance sheet accounts follow:

At December 31 2010 2009 2008
Assets
Cash $ 36,229 $ 42,780 $ 44,562
Accounts receivable, net 106,073 76,377 57,087
Merchandise inventory 137,408 98,929 62,038
Prepaid expenses 11,548 11,003 4,903
Plant assets, net 335,317 311,062 272,710
Total assets $ 626,575 $ 540,151 $ 441,300
Liabilities and Equity
Accounts payable $ 157,577 $ 94,024 $57,087
Long-term notes payable secured by mortgages on plant assets 116,618 127,962 99,478
Common stock, $10 par value 163,500 163,500 163,500
Retained earnings 188,880 154,665 121,235
Total liabilities and equity $ 626,575 $ 540,151 $ 441,300

The company's income statements for the years ended December 31, 2010 and 2009, follow:

For Year Ended December 31 2010 2009
Sales $ 685,000 $ 557,000
Cost of goods sold $ 417,850 $ 356,265
Other operating expenses 207,282 141,971
Interest expense 8,175 8,960
Income taxes 12,900 12,450
Total costs and expenses 646,207 519,646
Net income $ 38,793 $ 37,354
Earnings per share $ 2.37 $ 2.28

Compare the company's long-term risk and capital structure positions at the end of 2010 and 2009 by computing the following ratios.


(c) Times interest earned.
Answer #1

Timed interest earned = Earning before interest and tax/Interest expense

2010 = (38793+12900+8175)/8175 = 7.32 Times

2009 = (37354+12450+8960)/8960 = 6.56 Times